If A and B are two square matrices such that B = A inverse then A*B =_________^^

Grade:12

2 Answers

Arun
5 years ago
Dear student
 
here A * B = A * A-1 = I
 
hope it helps
In case of any query, please feel free to ask
 
Regards
Arun
shahid afrid
5 years ago
If A & B are 2 square matrices such that B=A-1   
then, A*B = A*(A-1)  =  I                 (Capital i ) .
 
Hope your question is answered.
For further clarification of doubts, feel free to ask.
Happy to help! :)